# Algorithm Generation via Creative Ideation

### v1

@@ Abstract
Designing system algorithms remains challenging, where the discontinuous nature of the solution space often forces system engineers to rely on generic heuristics at the expense of performance. We study whether LLMs can practically drive algorithm generation, and find that they are biased towards well-known generic designs, rather than making the creative leaps needed to navigate the discontinuous solution space. To address this limitation, we introduce MetaMuse, a framework for creative ideation built on three self-reflection principles: (1) quantifying solution diversity and usefulness in measurable performance space, rather than abstract idea space, (2) steering ideation through external stimuli, rather than internal randomness, and (3) constructing executable solutions using waypoint reasoning, rather than free-form chain-of-thought. Extensive evaluation shows that MetaMuse can generate high-performing solutions for two critical problems at a global cloud provider: cache replacement (reducing cache misses by up to 35.76%) and online bin packing (reducing bin usage by up to 30.93%).
@@ 1 Introduction
Designing system algorithms continues to be a central challenge in computing systems. Traditionally, the development of such algorithms has been a manual and labor-intensive process. Our experience at a global cloud provider indicates that even seemingly simple algorithms used in production — such as cache replacement for data storage or bin packing for job scheduling — can require tens of thousands of engineering hours to design. As a result, practitioners often resort to generic heuristics from the literature, e.g., least-recently used (LRU) for cache replacement and first-fit for bin packing, which frequently result in suboptimal performance.
@@
Due to this discontinuity, we approach the algorithm generation task from a different angle, and frame it as a sampling process in the solution space. LLM attempts to generate distinct solutions at each step. This generative process represents a sequence of leaps in discontinuous solution space (Bubeck et al., 2023), which we formulate as creative ideation for LLMs. In fact, the systems community has long hypothesized algorithm design as a discovery process of ideas (Kant, 1985).
To study the algorithm generation task, we focus on high-impact problems at a global cloud provider: cache replacement and online bin packing. Our initial attempts of repeatedly sampling GPT-4o, Llama3.3-70B, and DeepSeek-V3 show that LLMs are fundamentally hindered by availability bias (Tversky & Kahneman, 1973) — LLMs are trained to output the most likely sequence of words, according to training datasets. As a result, solutions tend to cluster around well-known heuristics in the literature, e.g., least-recently (LRU) and least-frequently used (LFU) for caching. Furthermore, we find that this bias cannot be practically addressed through LLM hyperparameters like temperature (Ackley et al., 1985).
The key to creative ideation is exploiting knowledge that LLMs assume to be probabilistically irrelevant to the given problem. What is missing is a self-reflection process, which thinks how to generate subsequent solutions by inspecting what solutions have been generated so far. In realizing such a self-reflection framework, MetaMuse, we see three model-agnostic principles surfacing to best guide this self-reflection. First, evaluating diversity of generated solutions should be grounded in the measurable feedback space (e.g., simulation performance of system algorithms), rather than the abstract idea space (Sankar & Sen, 2024). Second, steering the ideation is achieved through external stimuli (e.g., keywords), rather than internal randomness (Honghua Chen and Nai Ding, 2023). Third, developing executable solutions from external stimuli takes structured checkpoint-based steps, or waypoint reasoning, rather than free-form chain-of-thought (Wei et al., 2023; Mehrotra et al., 2024).
@@
Empirical analysis of LLMs’ fundamental limitations in the algorithm generation task (§2) — shedding light on the impracticality of relying on LLMs’ internal randomness for ideation, based on two critical problems at a global cloud provider: cache replacement and online bin packing.
@@
Principled framework for creative ideation, MetaMuse (§3) — systematically combining three self-reflection principles, to guide the discontinuous leaps necessary to overcome LLMs’ bias.
@@
Practical evaluation of the algorithm generation task (§4) — showing that (1) MetaMuse generates better cache replacement and bin packing solutions, outperforming LLM-based baselines (by up to 9.89% fewer cache misses, and up to 21.06% less bin usage) and human heuristics (by up to 35.76% fewer cache misses, and up to 30.93% less bin usage); (2) MetaMuse has up to 1.78 more distinct cache replacement solutions and 1.80 more distinct bin packing solutions, than LLM-based baselines; (3) MetaMuse has a low per-solution cost, up to 4.93 cents with GPT-4o.
@@ 2.1 System Algorithm Design
System algorithms define how computing systems behave. They are typically designed to optimize some performance objectives (e.g., cache hit ratio), for some scenarios or workloads (e.g., web servers). The difficulty of designing such algorithms arises from the nature of their solution space. It is discontinuous, where even a small change in algorithm design (e.g., the use of data structures) can lead to sharp and non-linear changes in performance. In addition, the discontinuity does not provide sufficiently predictable patterns or smooth landscape, to guide the search. This is a significant departure from prior efforts on auto-tuning system config parameters (Alipourfard et al., 2017; Cortez et al., 2017; Liang et al., 2020), which can be formulated as numerical optimization in most cases.
To this end, our work explores the use of LLMs to design heuristic algorithms, i.e., the algorithm generation task. Due to the discontinuous solution space, we approach this task as a sampling process in this space. At each step, LLMs attempt to generate distinct solutions. Conceptually, solving such a discontinuous task requires certain “Eureka” ideas that constitute leaps in the progress towards the final solution (Bubeck et al., 2023). Here, we refer to these leaps as creative ideation.
@@ 2.2 Creative Ideation
For the algorithm generation task, creative ideation can be formulated as the following iterative process. Each iteration takes in problem statements of the algorithm, along with the set of functions to be implemented. For example, most caching algorithms can be abstracted into insert and evict functions (Yang et al., 2020). We can also take in metadata and feedback of previous solutions, as optional inputs. At iteration , the output is an executable cache solution () with all functions coded. We can optionally provide feedback to iteration , by measuring hit ratio of previous solutions (, …, ) on user-given workload trace, in an environment such as simulators. After cache solutions, we select the best-performing one from (, …, ).
@@ 2.3 LLM Limitations in Creative Ideation
Our work is motivated by the observation, where LLMs’ creative ideation capability is fundamentally limited by the same mechanism that enables it — the next-token prediction. In cognitive terms, responses are influenced by what one has frequently or recently seen, i.e., availability bias. For LLMs, this bias arises from the data frequency distribution in the training dataset.
Impacts on algorithm generation.
Due to this bias, LLMs tend to generate certain designs. We illustrate with cache replacement, using GPT-4o, Llama3.3-70B and DeepSeek-V3.
Figure 1 shows our attempts with Repeated Sampling (Brown et al., 2024; Snell et al., 2024). At iteration , each LLM is prompted to create a new design , with the default temperature of 1. Then, is implemented by GPT-4o in Python, and benchmarked in the simulator to measure hit ratios on 30 traces. Figure 1 clusters 1,200 solutions into centroids of human heuristics, according to the similarity of hit ratios. And, LLMs tend to design caches behaving similarly to well-known heuristics: least-recently used (LRU), least-frequently used (LFU), and first-in-first-out (FIFO). Another observation is that making Repeated Sampling aware of previous solutions does not mitigate this bias.
Is adjusting LLM hyperparameters all you need?
One such parameter is temperature (Ackley et al., 1985), which internally regulates randomness in the generative process. Temperature smooths the probability distribution of next-token candidates, as computed by the softmax of their logits. At higher temperatures, high probabilities are decreased, and low probabilities are increased.
Unfortunately, high temperatures only mitigate availability bias, without practically addressing it. The reason is that increasing temperature smooths the probability distribution, which is a monotonic transformation retaining the relative ranking of output token candidates. Even at extreme temperatures, the distribution would approach flat and result in incoherent LLM outputs.
@@ 2.4 Related Work
The first category is automatic heuristic design (Liu et al., 2024b; Ye et al., 2024; Zheng et al., 2025; Romera-Paredes et al., 2024; Novikov et al., 2025; Sharma, 2025). These efforts mostly rely on LLMs to evolve an initial population of candidate designs, through mutations and crossovers. This dependence on the initial population inherently constrains the reachable solution space. In addition, availability bias can impact how the population is evolved. Our case study highlights a bias towards frequently tweaking the cache algorithm’s scoring function, rather than leaping to other design dimensions. MetaMuse addresses these limitations by making leaps in the solution space.
There are also efforts exploring human-LLM collaborations to produce diverse outputs (Liu et al., 2024c; Wan et al., 2024; Vaccaro et al., 2024). However, they require human efforts and can be subject to human bias. Instead, MetaMuse posits that LLMs themselves are capable of thinking “outside the box” (Mehrotra et al., 2024), and proposes to combine three self-reflection principles to achieve this capability (§3).
@@ 3 MetaMuse
Towards effective creative ideation, we address availability bias through a principled framework of self-reflection, MetaMuse. The key is to enable LLMs to think “outside the box,” i.e., exploiting knowledge that otherwise seems probabilistically irrelevant to solving the user problem. Figure 2 shows the overview of an iteration, which outputs one executable solution. The first step is to evaluate the diversity of generated solutions so far (§3.1). This evaluation is based on the embedding computed from each solution’s feedback (e.g., cache hit ratios in the figure). The second step is to steer ideation, by selecting a set of stimuli (e.g., the “flower” keyword in the figure) based on the diversity evaluation. §3.2 describes two stimuli selection strategies in the current implementation: RSDict and RSDict-SF. Finally, MetaMuse develops stimuli into executable solutions (§3.3), through four waypoints of reasoning: property extraction, problem mapping, solution formulation, and code generation.
@@ 3.1 Evaluating Diversity
MetaMuse grounds diversity evaluation in feedback space, rather than idea space. In our case study, the former refers to performance measurements such as cache hit ratio, and the latter refers to the semantic embedding of algorithm descriptions or code implementations. Therefore, MetaMuse represents each solution with a feedback embedding, which is a vector of feedback measurements on different workload traces.
This principle is motivated by the observation that semantic differences do not necessarily imply distinct solutions. Considering a case where GPT-4o generated functionally equivalent LFU-based solutions twice in different descriptions: “The cache evicts the object that is least frequently accessed,” and “The cache evicts the object with the lowest priority. Upon a cache hit, the object’s priority is incremented.” Code semantics exhibit similar shortcomings, where a solution implemented with a linked list or a priority queue can have the same hit ratio.
Furthermore, feedback embedding has two desirable properties. First, since each dimension is a quantitative metric conveying magnitude, the Euclidean distance of two embeddings is a direct indication of how different corresponding solutions are. For the LFU case above, those two solutions would be considered equivalent because their distance is 0. Second, since feedback metrics typically pre-define a fixed range (e.g., 0–100% for cache hit ratio), the -dimensional embedding space is inherently structured and bounded. This allows us to compute the steering direction (§3.2).
@@ 3.2 Steering with External Stimuli
The objective of steering is to guide LLMs to generate solutions, which targets regions of the feedback embedding space. Rather than relying on LLMs’ internal randomness, MetaMuse uses external stimuli as the starting point for ideation. Stimuli should be unbiased to the problem, forcing LLMs to associate with knowledge that seems probabilistically irrelevant. One domain-agnostic instantiation of stimuli is keywords from an English dictionary.
Realizing this principle of steering has practical challenges. The first is the source of stimuli. In the case of keywords, a large dictionary offers a broader keyword coverage, but it also contains many keywords of little value. An example is technical jargon, e.g., “prolamin” in biology. In our case study, if LLMs are forced to associate these keywords, they would frequently use them as variable names, rather than inspiration to ideate new solutions. The second challenge is selecting stimuli at each iteration. §3.2.1 next presents selection strategies in our current implementation.
@@ 3.2.1 Stimuli Selection Strategies
RSDict.
RSDict is a stateless strategy that always randomly selects keywords from the dictionary. Since RSDict does not rely on evaluation results of previous solutions, it can steer scenarios where evaluation is infeasible or costly (e.g., human judges).
RSDict-SF.
Unlike RSDict, RSDict-SF considers the feedback embeddings of previous solutions, to compute the steering direction at each iteration. Steering direction is specified by a feedback embedding, which serves as the objective that RSDict-SF should aim to achieve in subsequent solutions. In our case study, steering direction enables both exploration (for diversity) and exploitation (for usefulness). For the former, the steering direction should be the farthest point from all previous solutions, i.e., the feedback embedding that has the greatest Euclidean distance from all other embeddings. For the latter, we set all dimensions of the target embedding vector to be a high value (e.g., 100% cache hit ratio).
Given an embedding of steering direction, RSDict-SF computes the set of stimuli that would likely develop into solutions close to it. Our current implementation formulates this step as a prediction problem, and solves it with Gaussian Process Regression (GPR) models. Given the sum of semantic embeddings of stimuli, we use GPR models (, …, ), to predict each of the dimensions in embedding. Having (, …, ) allows RSDict-SF to predict the expected feedback embedding of any set of stimuli. GPR models are fitted with all previous solutions at each iteration. We use the dot-product kernel for GPR models. Dot-product kernel can capture the pairwise semantic similarity between all observations across two solutions and is invariant to their ordering. For two solutions and , their similarity captured by is
where denotes the observations derived from ’s stimuli, and similar for .
We highlight two implementation details. First, predicting for all possible sets of stimuli is not feasible in practice, and a dictionary with 3,000 common English words would already have 3,000s sets. To this end, MetaMuse exercises the power-of-two random choices (Mitzenmacher et al., 2001). It randomizes two sets of stimuli from the dictionary, and predicts their expected feedback embeddings. Then, MetaMuse selects the one closest to the target embedding. Second, RSDict-SF bootstraps GPR models, with solutions generated by RSDict.

### v2

# MetaMuse: Algorithm Generation via Creative Ideation

@@ Abstract
Designing system algorithms remains challenging, where the discontinuous nature of the solution space often forces system engineers to rely on generic heuristics at the expense of performance. We study whether LLMs can practically drive algorithm generation, and find that they are biased towards well-known generic designs, rather than making the creative leaps needed to navigate the discontinuous solution space. To address this limitation, we introduce MetaMuse, a framework for creative ideation built on three self-reflection principles: (1) quantifying solution diversity and usefulness in measurable performance space, rather than abstract idea space, (2) steering ideation through external stimuli, rather than internal randomness, and (3) constructing executable solutions using waypoint reasoning, rather than free-form chain-of-thought. Considering two critical online problems at a global cloud provider, extensive evaluations show that MetaMuse can generate high-performing solutions: it reduces cache misses by up to 35.76% in cache replacement and reduces bin usage by up to 30.93% in online bin packing.
@@ 1 Introduction
Designing system algorithms continues to be a central challenge in computing systems. Traditionally, the development of such algorithms has been a manual and labor-intensive process. Our experience at a global cloud provider indicates that even seemingly simple algorithms used in production—such as cache replacement for data storage or bin packing for job scheduling—can require tens of thousands of engineering hours to design. As a result, practitioners often resort to generic heuristics from the literature, e.g., least-recently used (LRU) and least-frequently used (LFU) for cache replacement, and first-fit for bin packing, which frequently result in suboptimal performance.
@@
Due to this discontinuity, we approach the algorithm generation task from a different angle, framing it as a sampling process in the solution space where the LLM attempts to generate distinct solutions at each step. This generative process represents a sequence of leaps across the discontinuous solution space (Bubeck et al., 2023), which we formulate as creative ideation for LLMs. In fact, the systems community has long hypothesized algorithm design as a discovery process of ideas (Kant, 1985).
To study the algorithm generation task, we focus on high-impact problems at a global cloud provider: cache replacement and online bin packing. Our initial attempts of repeatedly sampling GPT-4o, Llama3.3-70B, and DeepSeek-V3 show that LLMs are fundamentally hindered by availability bias (Tversky & Kahneman, 1973)—LLMs are trained to output the most likely sequence of words, according to training datasets. As a result, solutions tend to cluster around well-known heuristics in the literature, e.g., LRU and LFU for caching. Furthermore, we find that this bias cannot be practically addressed through LLM hyperparameters like temperature (Ackley et al., 1985).
The key to creative ideation is exploiting knowledge that LLMs assume to be probabilistically irrelevant to the given problem. What is missing is a self-reflection process, which thinks how to generate subsequent solutions by inspecting what solutions have been generated so far. In realizing such a self-reflection framework, MetaMuse, we see three model-agnostic principles surfacing to best guide this self-reflection. First, evaluating diversity of generated solutions should be grounded in the measurable feedback space (e.g., simulation performance of system algorithms), rather than the abstract idea space (Sankar & Sen, 2024). Second, steering the ideation is achieved through external stimuli (e.g., keywords), rather than internal randomness (Chen & Ding, 2023). Third, developing executable solutions from external stimuli takes structured checkpoint-based steps, or waypoint reasoning, rather than free-form chain-of-thought (Wei et al., 2023; Mehrotra et al., 2024).
@@
Empirical analysis of LLMs’ fundamental limitations in the algorithm generation task (§2). We shed light on the impracticality of relying on LLMs’ internal randomness for ideation, based on two online problems at a global cloud provider: cache replacement and online bin packing.
@@
Principled framework for creative ideation, MetaMuse (§3). We systematically combine three self-reflection principles to guide the discontinuous leaps necessary to overcome LLMs’ bias.
@@
Practical evaluation of the online algorithm generation task (§4). We show that (1) MetaMuse generates better cache replacement and bin packing solutions, outperforming LLM-based baselines (by up to 9.89% fewer cache misses, and up to 21.06% less bin usage) and human heuristics (by up to 35.76% fewer cache misses, and up to 30.93% less bin usage); (2) MetaMuse has up to 1.78 more distinct cache replacement solutions and 1.80 more distinct bin packing solutions, than LLM-based baselines; (3) MetaMuse has a low per-solution cost, up to 2.16 cents with GPT-4o.
@@ 2.1 System Algorithm Design
System algorithms define how computing systems behave. They are typically designed to optimize some performance objectives (e.g., cache hit ratio), for some scenarios or workloads (e.g., web servers). The difficulty of designing such algorithms arises from the nature of their solution space. It is discontinuous, where even a small change in algorithm design (e.g., the use of data structures) can lead to sharp and non-linear changes in performance. In addition, the discontinuity does not provide sufficiently predictable patterns or a smooth landscape to guide the search. This is a significant departure from prior efforts on auto-tuning system config parameters (Alipourfard et al., 2017; Cortez et al., 2017; Liang et al., 2020), which can be formulated as numerical optimization in most cases.
To this end, our work explores the use of LLMs to design heuristic algorithms, i.e., the algorithm generation task. Due to the discontinuous solution space, we formulate this task as a sampling process over the solution space, in which LLMs aim to generate distinct solutions at each step. Conceptually, solving such a discontinuous task requires certain “Eureka” ideas that constitute leaps towards the final solution (Bubeck et al., 2023). Here, we refer to these leaps as creative ideation.
@@ 2.2 Creative Ideation
For the algorithm generation task, creative ideation can be formulated as the following iterative process. Each iteration takes in the problem statement of the algorithm, along with the set of functions to be implemented. For example, most caching algorithms can be abstracted into insert and evict functions (Yang et al., 2020). Optionally, we can also incorporate metadata and feedback from previously generated solutions. At iteration , the output is an executable cache solution, , with all functions coded. We can optionally provide feedback to iteration , by evaluating the hit ratios of previous solutions (, …, ), on a user-given workload trace in an environment such as simulators. After cache solutions, we select the best-performing one from (, …, ).
@@ 2.3 LLM Limitations in Creative Ideation
Our work is motivated by the observation that the creative ideation capability of LLMs is fundamentally constrained by the very mechanism that enables it—next-token prediction. In cognitive terms, responses are shaped by what has been frequently or recently seen, a phenomenon known as availability bias. For LLMs, this bias stems from the frequency distribution of data seen in training.
Impacts on algorithm generation. Due to this bias, LLMs tend to generate certain classical designs. We illustrate this phenomenon in the context of cache replacement, using GPT-4o (version: 1120), Llama3.3-70B (version: Instruct), and DeepSeek-V3 (version: 0324).
Figure 1 shows our attempts with Repeated Sampling (Brown et al., 2024; Snell et al., 2024). At iteration , each LLM is prompted to generate a new algorithm design using a temperature of 1. To isolate creativity from coding ability, is then implemented in Python by the same model, GPT-4o. The resulting implementation is evaluated in a simulator to measure hit ratios on 30 synthetic traces. We examine two variants of Repeated Sampling that differ in whether the prompt at each iteration includes all previously generated solutions (i.e., solution descriptions and implementations). Figure 1 clusters 1,200 solutions by assigning them to the nearest centroids of established human heuristics, based on the similarity of their hit-ratio vectors. We observe that LLM-generated solutions concentrate around well-known heuristics, such as LRU, LFU, and first-in-first-out (FIFO). Moreover, making Repeated Sampling aware of previous solutions does not mitigate this bias.
Even if LLMs are instructed to generate new solutions by mutating previous solutions (Novikov et al., 2025; Sharma, 2025), we observe a bias towards tweaking the solution’s scoring function. LLMs tend to ignore other design dimensions such as data structure, hierarchical architecture, etc.
Is adjusting LLM hyperparameters all you need? One such parameter is temperature (Ackley et al., 1985), which internally regulates randomness in the generative process. Temperature smooths the probability distribution of next-token candidates, as computed by the softmax of their logits. At higher temperatures, high probabilities are decreased, and low probabilities are increased.
Unfortunately, high temperatures only marginally mitigate availability bias without fundamentally addressing it. The reason is that increasing temperature smooths the probability distribution, which is a monotonic transformation retaining the relative ranking of output token candidates. Even at extreme temperatures, the distribution would approach uniformity, resulting in incoherent LLM outputs.
@@ 2.4 Related Work
The first category is automatic heuristic design (Liu et al., 2024b; Ye et al., 2024; Zheng et al., 2025; Romera-Paredes et al., 2024; Novikov et al., 2025; Sharma, 2025; Dwivedula et al., 2025; Lange et al., 2025; Cheng et al., 2025). These efforts mostly rely on LLMs to evolve an initial population of candidate designs through mutations and crossovers. This dependence on the initial population inherently constrains the reachable solution space. In addition, availability bias can influence how the population evolves. Our case study reveals a bias towards frequently tweaking the cache algorithm’s scoring function, rather than leaping to other design dimensions. MetaMuse addresses these limitations by making leaps in the solution space.
There are also efforts exploring human-LLM collaborations to produce diverse outputs (Liu et al., 2024c; Wan et al., 2024; Vaccaro et al., 2024). However, they require human involvement and can be subject to human biases. In contrast, MetaMuse posits that LLMs themselves are capable of thinking “outside the box” (Mehrotra et al., 2024), and proposes combining three self-reflection principles to achieve this capability (§3).
@@ 3 MetaMuse
In this work, we introduce MetaMuse, a creative ideation framework that addresses availability bias through principled self-reflection. The key is to enable LLMs to think “outside the box,” leveraging knowledge that otherwise seems probabilistically irrelevant to the user’s problem.
Figure 2 illustrates a single iteration of MetaMuse, which outputs one executable solution. The first step evaluates the diversity of solutions generated so far (§3.1) based on embedding vectors constructed from each solution’s feedback (e.g., cache hit ratios, as shown in the figure). The second step steers ideation by selecting a set of stimuli (e.g., the keyword “flower” in the figure). §3.2 describes two stimuli selection strategies in the current implementation: RSDict and RSDict-SF. Finally, MetaMuse develops the stimuli into executable solutions (§3.3), through four waypoints of reasoning: property extraction, problem mapping, solution formulation, and code generation.
@@ 3.1 Evaluating Diversity
MetaMuse grounds diversity evaluation in feedback space rather than idea space. In our case study, the former refers to performance measurements such as cache hit ratios, whereas the latter refers to semantic embeddings of algorithm descriptions or code implementations. MetaMuse represents each solution using a feedback embedding, which is a vector containing performance measurements across workload traces. Ideally, these traces should be diverse, and one approach is to randomly synthesize them using tools from the systems community (e.g., libCacheSim (Yang et al., 2020) for cache workloads).
This principle is motivated by the observation that semantic differences do not necessarily imply distinct solutions. Considering a case where GPT-4o generated functionally equivalent LFU-based solutions twice from different design descriptions: “The cache evicts the object that is least frequently accessed,” and “The cache evicts the object with the lowest priority. Upon a cache hit, the object’s priority is incremented.” Code semantics exhibit similar shortcomings, where a solution implemented with a linked list or a priority queue can yield the same hit ratio.
Furthermore, feedback embeddings have two desirable properties. First, because each dimension is a quantitative metric conveying magnitude, the Euclidean distance between two embeddings is a direct indication of how different the corresponding solutions are. In the LFU example above, those two solutions would be considered equivalent as their distance is 0. Second, feedback metrics typically lie within a fixed range, either by definition or through normalization (e.g., 0–100% for cache hit ratio). As a result, the -dimensional embedding space is inherently structured and bounded, allowing us to compute a well-defined steering direction (§3.2).
@@ 3.2 Steering with External Stimuli
The objective of steering is to guide LLMs to generate solutions, targeting regions of the feedback embedding space. Rather than relying on the model’s internal randomness, MetaMuse introduces external stimuli as the starting point for ideation. These stimuli may be unrelated to the problem domain, forcing LLMs to associate with knowledge that appears probabilistically irrelevant. One domain-agnostic instantiation of such stimuli is a set of keywords from an English dictionary.
Realizing this principle of steering presents practical challenges. The first is the source of stimuli. In the case of keywords, while a large dictionary provides broad keyword coverage, it also includes many keywords of little value. An example is technical jargon, e.g., “prolamin” in biology. In our case study, if LLMs are forced to associate these keywords, they would frequently use them as variable names, rather than as inspiration to ideate new solutions. The second challenge is selecting stimuli at each iteration. §3.2.1 next presents selection strategies in our current implementation.
@@ 3.2.1 Stimuli Selection Strategies
RSDict. RSDict is a stateless strategy that randomly selects keywords from the dictionary. Since RSDict does not rely on the evaluation results of previous solutions, it is applicable in settings where evaluation is infeasible or costly (e.g., when human judges are required).
RSDict-SF. Based on RSDict, RSDict-SF further relies on self-feedback, leveraging the feedback embeddings of prior solutions to compute a steering direction. The steering direction is specified as a target feedback embedding, which serves as the objective that RSDict-SF should aim to achieve in subsequent solutions. In our case study, this mechanism supports both exploration (for diversity) and exploitation (for usefulness). For exploration, the steering direction is defined as the point in the embedding space that is farthest from all previous solutions, i.e., the embedding with the greatest Euclidean distance from the existing set. For exploitation, we define the target embedding by setting all dimensions to a high value (e.g., 100% cache hit ratios across all traces).
Given a steering-direction embedding, RSDict-SF computes a set of stimuli that are likely to develop into solutions close to the target. Our current implementation formulates this step as a prediction problem and solves it with Gaussian Process Regression (GPR). Specifically, given stimuli, we use their corresponding observations to predict the solution’s feedback embedding dimension by dimension, without generating the solution itself. Each observation is a natural-language sentence generated by LLMs after mapping one stimulus to the problem domain; we describe this process in detail in §3.3. We first compute the semantic embeddings (e.g., by SBERT Reimers et al. (2019)) of the observations, and use GPR models, (, …, ), to predict each of the dimensions in feedback embedding. These models allow RSDict-SF to predict the expected feedback embedding for any set of stimuli without generating the solution. The GPR models are refitted at each iteration with all previous solutions. We adopt a dot-product kernel for the GPR models. This kernel captures pairwise semantic similarity between all observations across two solutions and is invariant to their ordering. For two solutions and , their similarity under kernel is
where denotes the -th observation generated from the -th stimulus in the set of stimuli for , and similarly for . denotes the semantic embedding of an observation.
We highlight two implementation details. First, predicting for all possible sets of stimuli is not feasible in practice, and a dictionary with 3,000 common English words would already have 3,000s sets. To this end, MetaMuse exercises the power-of-two random choices (Mitzenmacher et al., 2001). It randomizes two sets of stimuli from the dictionary, and predicts their expected feedback embeddings. Then, MetaMuse selects the one closest to the target embedding. Second, RSDict-SF is bootstrapped, with solutions generated by RSDict, in order to start training GPR models.